# Dynamically Anchored Prompting for Task-Imbalanced Continual Learning

### v1

@@ Abstract
Existing continual learning literature relies heavily on a strong assumption that tasks arrive with a balanced data stream, which is often unrealistic in real-world applications. In this work, we explore task-imbalanced continual learning (TICL) scenarios where the distribution of task data is non-uniform across the whole learning process. We find that imbalanced tasks significantly challenge the capability of models to control the trade-off between stability and plasticity from the perspective of recent prompt-based continual learning methods. On top of the above finding, we propose Dynamically Anchored Prompting (DAP), a prompt-based method that only maintains a single general prompt to adapt to the shifts within a task stream dynamically. This general prompt is regularized in the prompt space with two specifically designed prompt anchors, called boosting anchor and stabilizing anchor, to balance stability and plasticity in TICL. Remarkably, DAP achieves this balance by only storing a prompt across the data stream, therefore offering a substantial advantage in rehearsal-free CL. Extensive experiments demonstrate that the proposed DAP results in 4.5% to 15% absolute improvements over state-of-the-art methods on benchmarks under task-imbalanced settings. Our code is available at https://github.com/chenxing6666/DAP
@@ 1 Introduction
While descending and ascending TICL are two extreme cases, the shuffled TICL can be regarded as the general form of TICL.
To show the challenges brought by the proposed scenarios, we conduct a quick experiment using DualPrompt Wang et al. (2022d), a typical prompt-based CL method, in Figure 1(b).
It can be observed that in the descending TICL case, the model initially learns well on data-rich tasks but rapidly declines when training becomes scarce. This indicates the poor plasticity of the model in adapting to new tasks with fewer samples. On the contrary, in the ascending TICL case, the model initially struggles with learning data-scarce tasks and remains poor in the following tasks. This is due to the severe forgetting caused by the following data-rich tasks (see Figure. 2 for more analysis). These results reveal a more realistic plasticity and stability dilemma: given an incoming task, the model learns poorly when its training data is scarce and forgets rapidly when there is abundant data. Undoubtedly, both learning and forgetting become more challenging in TICL. Therefore, there is a clear need to design a specific method to address this issue.
@@ 2.3 Imbalanced Continual Learning
It is worth noting that these studies are primarily rehearsal-based methods. This underscores the significant challenge of storing samples where class imbalance is inherent, as data selection itself is imbalanced. Furthermore, our approach to TICL diverges fundamentally from these existing works. While they concentrate on addressing imbalances within samples or classes, our focus is on the imbalance across different tasks.
@@ 3.1 Preliminaries
Our CL protocol adopts the class-incremental CL setting. The training data is denoted as a sequence of $T$ tasks $\mathcal{D}=\{\mathcal{D}_{1},\ldots,\mathcal{D}_{T}\}$, where $\mathcal{D}_{t}=\{(x_{i}^{t},y_{i}^{t})\},i=1,...,N_{t}$ is sampled from a joint data distribution in the input and label space $X_{t}\times Y_{t}$ at task $t$, whose size (i.e., the number of samples in this task) is denoted as $N_{t}$. The target model is formulated as $f:X\rightarrow Y$, integrating a patch embedding layer $f_{p}$, and a backbone $f_{b}$ consisting of a stack of transformer encoder layers followed by a classifier, thus $f=f_{p}\circ f_{b}$. We employ a ViT-Base model, pre-trained on ImageNet, as the frozen feature extractor. In this class-incremental setting, similar to recent prompt-based CL methodsSmith et al. (2023); Wang et al. (2022d, e), task boundaries are clearly defined with no shared classes between them, and task identity is provided only during training.
@@ 3.2 Task-Imbalanced Continual Learning
In this paper, we formally study task-imbalanced continual learning (TICL). TICL is characterized by the distribution of task sizes, i.e., $N_{t}$, throughout the learning process. Different from the conventional task-balanced CL where $N_{t}$ is drawn from a uniform distribution, i.e., $N_{i}=N_{j},\forall i\neq j~{}\text{and}~{}i,j\in\{1,...,T\}$. Any distribution deviating from the uniform one introduces an inherent imbalance among tasks. By the long-tail nature of real-world data distributions, we generally assume that the task sizes also follow the long-tail distribution. Specifically, the task sizes follow $N_{I_{1}}>...>N_{I_{T}},~{}\forall I_{i}<I_{j}$, where $I_{1},...,I_{T}$ are the sorting index based on the task size by non-increasing order.
@@ 3.3 Case Study
Figure 2: Performance analysis of DualPrompt on different TICL cases with evaluation metrics. (a): $F$ shows the model stability to resist forgetting. (b): $P$ shows the model plasticity to learn new tasks.
@@
Figure 3: Overview of the proposed dynamic anchored prompting (DAP) for task-imbalanced continual learning. Task-imbalanced training data stream represents the sequential arrival of data from Task 0 to Task t. There are two phases for each task. In in-task phase 1, the initialized task-specific prompt learns knowledge related to the current task, serving as a boosting anchor. In in-task phase 2, the general prompt is trained, with the assistance of boosting anchor to ensure plasticity. Meanwhile, the centers of all previously learned task-specific prompts serve as stabilizing anchors, emphasizing stability. It’s worth noting that the boosting anchors of past tasks are not stored as the stabilizing anchor is updated in an online manner.
@@
There exists a stability-plasticity dilemma in TICL scenarios. Accordingly, we propose a method to effectively balance this dilemma.
@@ 4.1 Anchored Prompting
Stabilizing Anchor. To ensure model stability, the stabilizing anchor $\mathbf{p}_{s}$ is designed to prevent knowledge forgetting from past tasks by monitoring the learned task-specific prompts. To maintain the knowledge of all of the learned tasks, $\mathbf{p}_{s}$ is calculated by the weighted center of the boosting anchors of all learned tasks in the prompt space from task 1 to the current task $t$. The weight is associated with the inverse of the size of each task. To make the algorithm rehearsal-free, i.e., without storing any data or prompt, $\mathbf{p}_{s}$ can be updated in an online manner:
$\mathbf{p}_{s}\leftarrow\frac{1}{\sum_{i=1}^{t}N_{i}}\left(\frac{\mathbf{p}_{s}}{\sum_{i=1}^{t-1}N_{i}}+\frac{\mathbf{p}_{b}}{N_{t}}\right).$ (3)
@@ 4.2 Dynamic Stability-Plasticity Regularization
To address the remaining issue in the DAP framework, we introduce a strategy that enables the model to adjust its focus between plasticity and stability in the nature of TICL tasks. In the second in-task phase, we optimize the general prompt $\mathbf{p}_{g}$ by considering the balance between stability and plasticity. Different from $\mathbf{p}_{b}$ that is initialized in each new task, $\mathbf{p}_{g}$ is initialized in task 1 and updated through the whole CL process in order to achieve the ability to generalize well on all learned classes. Thus, we propose dynamic stability-plasticity factor $\lambda$ as the coefficient between two anchor alignments. The factor $\lambda$ modulates the balance between stability and plasticity by considering the size of the current task $t$ and the sizes of past tasks:
$\lambda=\frac{N_{t}-N_{\text{min}}}{N_{\text{max}}-N_{\text{min}}+\epsilon},$ (5)
where $N_{\text{min}}=\min_{i=0,...,t}N_{i}$ and $N_{\text{max}}=\max_{i=0,...,t}N_{i}$ are the minimum and maximum sizes of learned tasks, and $\epsilon$ is a small positive constant to prevent division by zero. It is basically the min-max normalization that measures the difference between the size of the current task and the minimum size of the learned tasks, which is then normalized into the range of $[0,1]$. As with many re-weighting techniques in long-tail learning Peng et al. (2023); Chen et al. (2023); Zhang et al. (2023b), the task size $N_{t}$ is an important indicator to represent the learning difficulty. It is relatively easier to learn from a task with a larger size. Accordingly, we use the factor $\lambda$ and $1-\lambda$ as the regularization coefficient to reflect this relationship. In the second in-task phase, the general prompt updated in the current task $t$ is given by the following loss function with the dynamic stability-plasticity factor $\lambda$:
$\mathcal{L}_{2}=$ $\sum_{i=1}^{N_{t}}l_{CE}(f_{b}([\mathbf{p}_{g},f_{p}(x_{i}^{t})]),y_{i}^{t})$
@@ 5.1 Implementation Details
Comparison Methods. In our experiments, we compare our Dynamically Anchored Prompting (DAP) with two groups of rehearsal-based methods. The first group includes classical approaches like PODNET Douillard et al. (2020) and BiC Wu et al. (2019). The second group comprises methods designed for long-tailed distributions such as EEIF++ Liu et al. (2022) and LUCIR++ Liu et al. (2022). Additionally, we include pretrained methods in our comparison, FineTune Khan et al. (2023), iCaRL Rebuffi et al. (2017), both these methods start from the same ImageNet pre-trained ViT-Base Dosovitskiy et al. (2020) model to ensure a fair comparison. Lastly, we compare DAP against the current state-of-the-art (SOTA) prompt-based methods, including L2P Wang et al. (2022e), DualPrompt Wang et al. (2022d), and CODA-Prompt Smith et al. (2023). These methods represent the latest advancements in prompt-based CL.
@@ 6 Conclusion
In this paper, we formally define task-imbalanced continual learning (TICL) and systematically study its three cases. We discovered that imbalanced tasks significantly deteriorate the performance of prompt-based CL methods because they raise a new challenge to consider the dilemma of stability and plasticity. To counteract this, we introduced Dynamically Anchored Prompting (DAP). DAP addresses the challenge by separating stability and plasticity with two prompts, one for stabilization and the other for plasticity, serving as anchors to guide the learning process of a general prompt. DAP improves the performance of baseline prompt-based TICL methods to set a new state-of-the-art.
7 Acknowledgements

### v2

@@ Abstract
Existing continual learning literature relies heavily on a strong assumption that tasks arrive with a balanced data stream, which is often unrealistic in real-world applications. In this work, we explore task-imbalanced continual learning (TICL) scenarios where the distribution of task data is non-uniform across the whole learning process. We find that imbalanced tasks significantly challenge the capability of models to control the trade-off between stability and plasticity from the perspective of recent prompt-based continual learning methods. On top of the above finding, we propose Dynamically Anchored Prompting (DAP), a prompt-based method that only maintains a single general prompt to adapt to the shifts within a task stream dynamically. This general prompt is regularized in the prompt space with two specifically designed prompt anchors, called boosting anchor and stabilizing anchor, to balance stability and plasticity in TICL. Remarkably, DAP achieves this balance by only storing a prompt across the data stream, therefore offering a substantial advantage in rehearsal-free CL. Extensive experiments demonstrate that the proposed DAP results in 4.5% to 15% absolute improvements over state-of-the-art methods on benchmarks under task-imbalanced settings. Our code is available at https://github.com/chenxing6666/DAP.
@@ 1 Introduction
While descending and ascending TICL are two extreme cases, shuffled TICL can be regarded as the general form of TICL.
To show the challenges brought by the proposed scenarios, we conduct a quick experiment using DualPrompt Wang et al. (2022d), a typical prompt-based CL method, in Figure 1(b). It can be observed that in the descending TICL case, the model initially learns well on data-rich tasks but rapidly declines when training becomes scarce. This indicates the poor plasticity of the model in adapting to new tasks with fewer samples. On the contrary, in the ascending TICL case, the model initially struggles with learning data-scarce tasks and remains poor in the following tasks. This is due to the severe forgetting caused by the following data-rich tasks (see Figure. 2 for more analysis). These results reveal a more realistic plasticity and stability dilemma: given an incoming task, the model learns poorly when its training data is scarce and forgets rapidly when there is abundant data. Undoubtedly, both learning and forgetting become more challenging in TICL. Therefore, there is a clear need to design a specific method to address this issue.
@@ 2.3 Imbalanced Continual Learning
It is noteworthy that these studies are primarily rehearsal-based methods. This underscores the significant challenge of storing samples where class imbalance is inherent, as data selection itself is imbalanced. Furthermore, our approach to TICL diverges fundamentally from these existing works. While they primarily concentrate on addressing imbalances within samples or classes, our focus lies in the imbalance across different tasks.
@@ 3.1 Preliminaries
Our CL protocol adopts the class-incremental CL setting. The training data is denoted as a sequence of $T$ tasks $\mathcal{D}=\{\mathcal{D}_{0},\ldots,\mathcal{D}_{T}\}$, where $\mathcal{D}_{t}=\{(x_{i}^{t},y_{i}^{t})\},i=0,...,N_{t}$ is sampled from a joint data distribution in the input and label space $X_{t}\times Y_{t}$ at task $t$, whose size (i.e., the number of samples in this task) is denoted as $N_{t}$. The target model is formulated as $f:X\rightarrow Y$, integrating a patch embedding layer $f_{p}$, and a backbone $f_{b}$ consisting of a stack of transformer encoder layers followed by a classifier, thus $f=f_{p}\circ f_{b}$. We employ a ViT-Base model, pre-trained on ImageNet, as the frozen feature extractor. In this class-incremental setting, similar to recent prompt-based CL methodsSmith et al. (2023); Wang et al. (2022d, e), task boundaries are clearly defined with no shared classes between them, and task identity is provided only during training.
@@ 3.2 Task-Imbalanced Continual Learning
In this paper, we formally study task-imbalanced continual learning (TICL). TICL is characterized by the distribution of task sizes, i.e., $N_{t}$, throughout the learning process. Different from the conventional task-balanced CL where $N_{t}$ is drawn from a uniform distribution, i.e., $N_{i}=N_{j},\forall i\neq j~{}\text{and}~{}i,j\in\{0,...,T\}$. Any distribution deviating from the uniform one introduces an inherent imbalance among tasks. By the long-tail nature of real-world data distributions, we generally assume that the task sizes also follow the long-tail distribution. Specifically, the task sizes follow $N_{I_{0}}>...>N_{I_{T}},~{}\forall I_{i}<I_{j}$, where $I_{0},...,I_{T}$ are the sorting index based on the task size by non-increasing order.
@@ 3.3 Case Study
Figure 2: Performance analysis of DualPrompt on different TICL cases with evaluation metrics. (a) $F$ shows the model stability to resist forgetting. (b) $P$ shows the model plasticity to learn new tasks.
@@
Figure 3: Overview of the proposed dynamic anchored prompting (DAP) for task-imbalanced continual learning. Task-imbalanced training data stream represents the sequential arrival of data from Task 0 to Task t. There are two phases for each task. In in-task phase 1, the initialized task-specific prompt learns knowledge related to the current task, serving as a boosting anchor. In in-task phase 2, the general prompt is trained, with the assistance of boosting anchor to ensure plasticity. Meanwhile, the centers of all previously learned task-specific prompts serve as stabilizing anchors, emphasizing stability. It is worth noting that the boosting anchors of past tasks are not stored as the stabilizing anchor is updated in an online manner.
@@
Accordingly, we propose a method to effectively balance the stability-plasticity dilemma in TICL scenarios.
@@ 4.1 Anchored Prompting
Stabilizing Anchor. To ensure model stability, the stabilizing anchor $\mathbf{p}_{s}$ is designed to prevent knowledge forgetting from past tasks by monitoring the learned task-specific prompts. To maintain the knowledge of all of the learned tasks, $\mathbf{p}_{s}$ is calculated by the weighted center of the boosting anchors of all learned tasks in the prompt space from task 0 to the current task $t$. The weight is associated with the inverse of the size of each task. To make the algorithm rehearsal-free, i.e., without storing any data or prompt, $\mathbf{p}_{s}$ can be updated in an online manner:
$\mathbf{p}_{s}\leftarrow\frac{1}{\sum_{i=0}^{t}\frac{1}{N_{i}}}\left({\sum_{i=0}^{t-1}\frac{\mathbf{p}_{s}}{N_{i}}}+\frac{\mathbf{p}_{b}}{N_{t}}\right).$ (3)
@@ 4.2 Dynamic Stability-Plasticity Regularization
To address the remaining issue in the DAP framework, we introduce a strategy that enables the model to adjust its focus between plasticity and stability in the nature of TICL tasks. In the second in-task phase, we optimize the general prompt $\mathbf{p}_{g}$ by considering the balance between stability and plasticity. Different from $\mathbf{p}_{b}$ that is initialized in each new task, $\mathbf{p}_{g}$ is initialized in task 0 and updated through the whole CL process in order to achieve the ability to generalize well on all learned classes. Thus, we propose dynamic stability-plasticity factor $\lambda$ as the coefficient between two anchor alignments. The factor $\lambda$ modulates the balance between stability and plasticity by considering the size of the current task $t$ and the sizes of past tasks:
$\lambda=\frac{N_{t}-N_{\min}}{N_{\max}-N_{\min}+\epsilon},$ (5)
where $N_{\min}=\min\{N_{i}\}$ and $N_{\max}=\max\{N_{i}\}$ are the minimum and maximum sizes of learned tasks for $i=0,...,t$, and $\epsilon$ is a small positive constant to prevent division by zero. It is basically the min-max normalization that measures the difference between the size of the current task and the minimum size of the learned tasks, which is then normalized into the range of $[0,1]$. As with many re-weighting techniques in long-tail learning Peng et al. (2023); Chen et al. (2023); Zhang et al. (2023b), the task size $N_{t}$ is an important indicator to represent the learning difficulty. It is relatively easier to learn from a task with a larger size. Accordingly, we use the factor $\lambda$ and $1-\lambda$ as the regularization coefficient to reflect this relationship. In the second in-task phase, the general prompt updated in the current task $t$ is given by the following loss function with the dynamic stability-plasticity factor $\lambda$:
$\mathcal{L}_{2}=$ $\sum_{i=0}^{N_{t}}l_{CE}(f_{b}([\mathbf{p}_{g},f_{p}(x_{i}^{t})]),y_{i}^{t})$
@@ 5.1 Implementation Details
Comparison Methods. In our experiments, we compare the proposed DAP with two groups of rehearsal-based methods. The first group includes classical approaches like PODNET Douillard et al. (2020) and BiC Wu et al. (2019). The second group comprises methods designed for long-tailed distributions such as EEIF++ Liu et al. (2022) and LUCIR++ Liu et al. (2022). Additionally, we include pretrained methods in our comparison, FineTune Khan et al. (2023), iCaRL Rebuffi et al. (2017), both these methods start from the same ImageNet pre-trained ViT-Base Dosovitskiy et al. (2020) model to ensure a fair comparison. Lastly, we compare DAP against the current state-of-the-art (SOTA) prompt-based methods, including L2P Wang et al. (2022e), DualPrompt Wang et al. (2022d), and CODA-Prompt Smith et al. (2023).
@@ 6 Conclusion
In this paper, we formally define task-imbalanced continual learning and systematically study its three cases. We discovered that imbalanced tasks significantly deteriorate the performance of prompt-based CL methods because they raise a new challenge to consider the dilemma of stability and plasticity. To counteract this, we introduced dynamically anchored prompting. DAP addresses the challenge by separating stability and plasticity with two prompts, one for stabilization and the other for plasticity, serving as anchors to guide the learning process of a general prompt. DAP improves the performance of baseline prompt-based TICL methods to set a new state-of-the-art.
Acknowledgements